# Sharper Guarantees for Learning Neural Network Classifiers with Gradient Methods

### v1

@@ Abstract
In this paper, we study the data-dependent convergence and generalization behavior of gradient methods for neural networks with smooth activation. Our first result is a novel bound on the excess risk of deep networks trained by the logistic loss, via an alogirthmic stability analysis. Compared to previous works, our results improve upon the shortcomings of the well-established Rademacher complexity-based bounds. Importantly, the bounds we derive in this paper are tighter, hold even for neural networks of small width, do not scale unfavorably with width, are algorithm-dependent, and consequently capture the role of initialization on the sample complexity of gradient descent for deep nets. Specialized to noiseless data separable with margin $\gamma$ by neural tangent kernel (NTK) features of a network of width $\Omega(\mathrm{poly}(\log(n)))$, we show the test-error rate to be $e^{O(L)}/{\gamma^{2}n}$, where $n$ is the training set size and $L$ denotes the number of hidden layers. This is an improvement in the test loss bound compared to previous works while maintaining the poly-logarithmic width conditions. We further investigate excess risk bounds for deep nets trained with noisy data, establishing that under a polynomial condition on the network width, gradient descent can achieve the optimal excess risk. Finally, we show that a large step-size significantly improves upon the NTK regime’s results in classifying the XOR distribution. In particular, we show for a one-hidden-layer neural network of constant width $m$ with quadratic activation and standard Gaussian initialization that mini-batch SGD with linear sample complexity and with a large step-size $\eta=m$ reaches the perfect test accuracy after only $\lceil\log(d)\rceil$ iterations, where $d$ is the data dimension.
@@ 1.1 Overview
Yet, even for the relatively simple setting of learning deep nets in the kernel regime, the existing generalization bounds are still sub-optimal. Moreover the boundaries of the kernel regime are still largely unknown and an active area of research [Liu et al., 2022, Banerjee et al., 2022, Telgarsky, 2023]. While the kernel regime can partially demonstrate the behavior of neural networks, the resulting guarantees often require large width, small step-size or large iteration and sample complexities. There is increasing evidence in recent years that for certain class of data distributions neural networks can overcome these limitations by using large step-sizes which allows the network’s parameters to move a long distance from initialization often leading to better sample and computational complexities [Damian et al., 2022, Ba et al., 2022, Glasgow, 2024].
In this work, we study the generalization and convergence behavior of gradient-based algorithms in neural nets with smooth activation functions for a wide class of data distributions. Our first result characterizes the test and train loss rates for classification problems under the condition that deviation from initialization is bounded depending on the network’s width. In particular, for $L$-hidden layer networks our results hold under $\|w^{\star}-w_{0}\|\lesssim m^{O(1/L)}$, allowing the network’s weights to move from initialization a distance increasing with $m$. This shows that kernel regime continues to hold for a wider range of setups than previous results for which the deviations are constrained to be constant in $m$. The key reason for this improvement is exploiting the objective’s Hessian structure in the gradient descent (GD) path. More importantly, using the Hessian information enables us to develop, for the first time, algorithm-dependent generalization bounds of deep neural networks. As it will be discussed throughout the paper, the bounds we derive via algorithmic-stability are tighter than previous relevant bounds in literature (Table 1).
@@
Our analysis is inspired by the recent progress in bounding the multi-layer networks’ Hessian spectral norm [Liu et al., 2020, Banerjee et al., 2022] as well as the progress in understanding the loss landscape of neural net classifiers during gradient descent iterates [Richards and Kuzborskij, 2021, Taheri and Thrampoulidis, 2024]. In contrast with previous works, we then specialize these results to a well-known NTK separability condition tailored for noiseless data and show that our results substantially improve the upper-bounds on the test error performance while still allowing the width to be small, specifically poly-logarithmic on sample size (Table 1). A more detailed comparison is deferred to Section 1.2. We also consider the case of noisy data distributions and show that deep models are consistent, i.e., they can achieve the optimal test loss in the presence of noise as sample size grows.
While these results improve upon the existing bounds within the NTK regime, we show in Section 2.2 that using large step-size can further improve both the computational and sample complexities. In particular, we show for the stylized setup of data following the XOR distribution, a two-layer neural network with quadratic activation reaches zero test error after only $\log(d)$ steps of mini-batch Stochastic Gradient Descent (SGD) with an aggressive step-size $\eta=m$. A comparison of our findings with the guarantees of the kernel regime with both GD (as per Corollary 1; see also [Ji and Telgarsky, 2020, Taheri and Thrampoulidis, 2024]) and one-pass SGD [Telgarsky, 2023, Chen et al., 2021] and the most relevant work in the feature learning regime [Glasgow, 2024] is summarized in Table 2. Below is a summary of our contributions.
@@ 1.2 Prior works
Some recent works have pointed out the limitations of the kernel regime in understanding the full power of neural networks [Abbe et al., 2022]. In particular, as in the kernel regime, the networks weights are bounded not to move significantly from initialization, the learned features are not considerably different from those learned at initialization. On the other hand, it is hypothesized that neural networks can learn the true underlying features of the data distribution if the network weights are allowed (by large step-sizes or avoiding early-stopping) to move a large distance from initialization. This phenomenon was first proved for specific regression tasks where the labels essentially only depend on a small number of features, such as when $y=g(Ux)$ for $U\in\mathbb{R}^{k\times d}$ where $k\ll d$ [Damian et al., 2022, Ba et al., 2022, Abbe et al., 2022, Cui et al., 2024], using one large gradient-descent step leading to superior sample complexities compared to the kernel regime. For classification tasks, some focus has been on the XOR distribution(a.k.a. parities) [Wei et al., 2019]. Recent works have studied the problem of learning the $d$-dimensional XOR distribution using neural networks in both NTK and feature learning settings [Barak et al., 2022, Telgarsky, 2023, Xu et al., 2024, Taheri and Thrampoulidis, 2024, Glasgow, 2024]. Specifically, it has been shown that under NTK with a sufficiently small step size, a polynomially wide network requires $d^{2}$ GD steps and $d^{2}$ sample size. Some studies have achieved linear sample complexity for learning XOR [Bai and Lee, 2019, Glasgow, 2024, Telgarsky, 2023]; but these methods involve more computational effort compared to our results. The work most related to ours is [Glasgow, 2024], which demonstrated that with a particular Gaussian initialization, a ReLU network requires $\mathrm{poly}(\log(d))$ large SGD steps and $\mathrm{poly}(\log(d))$ neurons to learn XOR with linear sample complexity. In contrast, we show that by using a quadratic activation, learning this distribution requires only $\log(d)$ steps with a constant-width network, while maintaining the same linear sample complexity.
@@ 2.1 Train and test loss bounds in deep nets
Consider $L$-layer neural network model with the model defined as $\Phi(w,x):=\frac{1}{\sqrt{m}}W_{L+1}^{\top}(\frac{1}{\sqrt{m}}\sigma(W_{L}^{\top}\cdots\frac{1}{\sqrt{m}}\sigma(W_{1}^{\top}x)\cdots)$ where $\sigma$ is a $1$-smooth and $1$-Lipschitz activation function such that $\sigma(0)=0$. Moreover, let $\beta_{L}$ be a constant that only depends on $L$, let all parameters of the network be initialized as i.i.d. standard Gaussian and assume the step-size satisfies the condition of the descent lemma. Fix $T$ and assume the target weights vector $w^{\star}\in\mathbb{R}^{p}$ that obtain small training loss such that
@@
Moreover, the expected generalization gap satisfies with high probability over initialization,
@@
where $G_{0}$ is the Lipschitz parameter of network at initialization i.e., $\|\nabla\Phi(w_{0},\cdot)\|\leq G_{0}$ and the expectation is over the data sampling.
@@ 2.1.1 Specializing to the NTK-separablity condition
In words, the above assumption implies that the features learned by the gradient at initialization can be linearly separated by some weights $w$. This assumption is commonly used in deep learning literature for studying classification tasks [Chen et al., 2021, Ji and Telgarsky, 2020].
@@
Consider the same setup as Theorem 2.1 and let Assumption 1 hold. Define constant $B>0$ that bounds the model’s output at initialization i.e., $\forall i\in[n]:|\Phi(w_{0},x_{i})|<B$. Assume the width is large enough such that $m\geq\beta_{L}^{2}\left(\frac{2B+\log(1/{\epsilon})}{\gamma}\right)^{6L+4}.$ Then there exists $w^{\star}\in\mathbb{R}^{p}$ such that $F(w^{\star})\leq{\epsilon}$ and $w^{\star}$ lies at a bounded Euclidean distance from initialization such that $\|w^{\star}-w_{0}\|\leq\frac{1}{\gamma}(2B+\log(1/{\epsilon}))$.
@@
under the condition that $m=\Omega(\mathrm{poly}(\log(n))\cdot\beta_{L}/\gamma^{6L+4})$. We remark the test error bound does not have any explicit dependence on $m$. This in itself is not surprising as one intuitively expects the bound not to scale unfavorably with width. Although recent works have derived width independent generalization bounds for two-layer networks [Ji and Telgarsky, 2020, Telgarsky, 2023, Taheri and Thrampoulidis, 2024], we are not aware of any prior work proving width-independent bounds for learning multi-layer networks with GD. As discussed earlier in the introduction, the bound derived in the closely related work [Chen et al., 2021] scales unfavorably with $m$ as it grows at the rate $\sqrt{m/n}$. In fact, the authors refer to deriving width-independent bounds as an open problem in [Chen et al., 2021, Sec 3.1].
@@ 2.2 Overcoming NTK limitations for XOR dataset
Next, we show the previous NTK guarantees can be improved by using large step-sizes. We consider the stylized set-up of learning the $d-$dimensional XOR data distribution with two layer neural networks with quadratic activation which takes the following form:
@@
Assume mini-batch SGD with batch-size $n\geq d\cdot\log^{14}(d)$ and step-size $\eta=m$. Then, after $T=\log(d)$ iterations the test accuracy satisfies with probability at least $1-e^{\log(m)-\log^{2}(d)}-2e^{-\frac{m}{16}}-o_{d}(1)$ over the randomness of initialization and data sampling,
@@ 3 Numerical Results
Experiments on learning the XOR distribution.
@@ B.2 Proof for Generalization loss
Lemma 5 (Lem. B.3 in [Lei and Ying, 2020]).
@@ E.4 Empirical Gradient for multiple GD steps
$w^{1}$ $=w^{0}-\eta\nabla\hat{F}(w^{0}),$
$\mathbb{E}_{x}[w^{1}]$ $=w^{0}-\eta\mathbb{E}_{x}[\nabla\hat{F}(w^{0})],$
$w^{1}$ $=\mathbb{E}_{x}[w^{1}]-\eta(\nabla\hat{F}(w^{0})-\mathbb{E}_{x}[\nabla\hat{F}(w^{0})]).$
@@
$w^{t+1}$ $=w^{t}-\eta\nabla\hat{F}(w^{t}),$
$\mathbb{E}_{x}[w^{t+1}]$ $=\mathbb{E}_{x}[w^{t}]-\eta\mathbb{E}_{x}[\nabla\hat{F}(\mathbb{E}_{x}[w^{t}])],$
$w^{t+1}$ $=\mathbb{E}_{x}[w^{t}]+(w^{t}-\mathbb{E}_{x}[w^{t}])-\eta(\nabla\hat{F}(w^{t})-\mathbb{E}_{x}[\nabla\hat{F}(\mathbb{E}_{x}[w^{t}])])-\eta\mathbb{E}_{x}[\nabla\hat{F}(\mathbb{E}_{x}[w^{t}])].$
@@
$w^{t+1}-\mathbb{E}_{x}[w^{t+1}]$ $=-\eta\sum_{\tau=0}^{t}\nabla\hat{F}(w^{\tau})-\mathbb{E}_{x}\left[\nabla\hat{F}(\mathbb{E}_{x}[w^{\tau}])\right]$
$=-\eta\sum_{\tau=0}^{t}\underbrace{\nabla\hat{F}(w^{\tau})-\nabla\hat{F}(\mathbb{E}_{x}[w^{\tau}])}_{\text{Term I}}+\underbrace{\nabla\hat{F}(\mathbb{E}_{x}[w^{\tau}])-\mathbb{E}_{x}[\nabla\hat{F}(\mathbb{E}_{x}[w^{\tau}])]}_{\text{Term II}}.$
@@ E.5 Term II:
$-\nabla\hat{F}_{i}(\mathbb{E}_{x}[w^{t}])=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w^{t}_{i}]\rangle x_{j}y_{j}$
@@
$-\nabla\hat{F}_{i}^{k}(\mathbb{E}_{x}[w^{t}])$ $=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w_{i}^{t}]\rangle x_{j}(k)x_{j}(1)x_{j}(2)$
@@
$|\nabla\hat{F}_{i}^{k}(\mathbb{E}_{x}[w^{t}])|$ $\lesssim\frac{\log(dT)\log(dm)}{nm}\left(\frac{2\sqrt{n}(|\alpha_{1}^{t}|+\alpha_{2}^{t})}{\sqrt{d}}+\sqrt{n}\log(d)\right)$
@@
$-\nabla\hat{F}_{i}^{1}(\mathbb{E}_{x}[w^{t}])$ $=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w_{i}^{t}]\rangle x_{j}(2)$
@@
$-\nabla\hat{F}_{i}^{2}(\mathbb{E}_{x}[w^{t}])=\frac{a_{i}}{m}\mathbb{E}_{x}[w_{i}^{t}(1)]+O(\log(dm)\log(dT)\frac{|\alpha_{1}^{t}|+\alpha_{2}^{t}+\sqrt{d}\log(d)}{m\sqrt{nd}}).$ (50)
@@
$\left\|\nabla\hat{F}_{i}(\mathbb{E}_{x}[w^{t}])-\mathbb{E}_{x}[\nabla F_{i}(\mathbb{E}_{x}[w^{t}])]\right\|\lesssim\log(dm)\log(dT)\frac{|\alpha_{1}^{t}|+\alpha_{2}^{t}+\sqrt{d}\log(d)}{m\sqrt{n}}.$
@@
$\forall i,t:\left\|\nabla\hat{F}_{i}(\mathbb{E}_{x}[w^{t}])-\mathbb{E}_{x}[\nabla\hat{F}_{i}(\mathbb{E}_{x}[w^{t}])]\right\|\lesssim\log(dm)\log(dT)\frac{|\alpha_{1}^{t}|+\alpha_{2}^{t}+\sqrt{d}\log(d)}{m\sqrt{n}},$

### v2

@@ Abstract
In this paper, we study the data-dependent convergence and generalization behavior of gradient methods for neural networks with smooth activation. Our first result is a novel bound on the excess risk of deep networks trained by the logistic loss, via an alogirthmic stability analysis. Compared to previous works, our results improve upon the shortcomings of the well-established Rademacher complexity-based bounds. Importantly, the bounds we derive in this paper are tighter, hold even for neural networks of small width, do not scale unfavorably with width, are algorithm-dependent, and consequently capture the role of initialization on the sample complexity of gradient descent for deep nets. Specialized to noiseless data separable with margin $\gamma$ by neural tangent kernel (NTK) features of a network of width $\Omega(\mathrm{poly}(\log(n)))$, we show the test-error rate to be $e^{O(L)}/{\gamma^{2}n}$, where $n$ is the training set size and $L$ denotes the number of hidden layers. This is an improvement in the test loss bound compared to previous works while maintaining the poly-logarithmic width conditions. We further investigate excess risk bounds for deep nets trained with noisy data, establishing that under a polynomial condition on the network width, gradient descent can achieve the optimal excess risk. Finally, we show that a large step-size significantly improves upon the NTK regime’s results in classifying the XOR distribution. In particular, we show for a one-hidden layer neural network of constant width $m$ with quadratic activation and standard Gaussian initialization that SGD with linear sample complexity and with a large step-size $\eta=m$ reaches the perfect test accuracy after only $\lceil\log(d)\rceil$ iterations, where $d$ is the data dimension.
@@ 1.1 Overview
Yet, even for the relatively simple setting of learning deep nets in the kernel regime, the existing generalization bounds are still suboptimal. Moreover, the boundaries of the kernel regime are still largely unknown and an active area of research [Liu et al., 2022, Banerjee et al., 2022, Telgarsky, 2023]. While the kernel regime can partially demonstrate the behavior of neural networks, the resulting guarantees often require large width, small step-size or large iteration and sample complexities. There is increasing evidence in recent years that for certain class of data distributions neural networks can overcome these limitations by using a large step-size which allows the network’s parameters to move a long distance from initialization, often leading to better sample and computational complexities [Damian et al., 2022, Ba et al., 2022].
In this work, we study the generalization and convergence behavior of gradient-based algorithms in neural nets with smooth activation functions for a wide class of data distributions. Our first result characterizes the test and train loss rates for classification problems under the condition that deviation from initialization is bounded depending on the network’s width. In particular, for $L$-hidden layer networks our results hold under $\|w^{\star}-w_{0}\|\lesssim m^{O(1/L)}$, allowing the network’s weights to move from initialization a distance increasing with $m$. This shows that the kernel regime continues to hold for a wider range of setups than previous results for which the deviations are restricted to be constant in $m$. The key reason for this improvement is exploiting the objective’s Hessian structure in the gradient-descent path. More importantly, using the Hessian information enables us to develop, for the first time, algorithm-dependent generalization bounds of deep neural networks. As will be discussed throughout the paper, the bounds we derive via algorithmic-stability are tighter than previous relevant bounds in the literature.
@@
We specialize these results to a well-known NTK separability condition tailored for noiseless data and show that our results substantially improve the prior results on the test error performance while still allowing the width to be small, specifically poly-logarithmic on sample size (Table 1). A more detailed comparison is deferred to Section 1.2. We also consider the case of noisy data distributions and show that deep models are consistent, i.e., they can achieve the optimal test loss in the presence of noise as the sample size grows.
While these results improve upon the existing bounds within the NTK regime, we show in Section 2.2 that using a large step-size can further improve both the computational and sample complexities. In particular, we show for the stylized setup of data following the XOR distribution, a two-layer neural network with quadratic activation reaches zero test error after only $\log(d)$ steps of SGD with an aggressive step-size $\eta=m$. A comparison of our findings with the guarantees of the kernel regime with both GD [Taheri and Thrampoulidis, 2024] and one-pass SGD [Telgarsky, 2023, Chen et al., 2021] and the most relevant work in the feature learning regime [Glasgow, 2024] is summarized in Table 2.
Below is a summary of our contributions.
@@ 1.2 Prior works
Some recent works have pointed out the limitations of the kernel regime in understanding the full power of neural networks [Abbe et al., 2022]. In particular, as in the kernel regime, the networks weights are bounded not to move significantly from initialization, the learned features are not considerably different from those learned at initialization. On the other hand, it is hypothesized that neural networks can learn the true underlying features of the data distribution if the network weights are allowed (by large step-sizes or avoiding early-stopping) to move a large distance from initialization. This phenomenon was first proved for specific regression tasks where the labels essentially only depend on a small number of features, such as when $y=g(Ux)$ for $U\in\mathbb{R}^{k\times d}$ where $k\ll d$ [Damian et al., 2022, Ba et al., 2022, Abbe et al., 2022, Cui et al., 2024], by one large SGD step leading to superior sample complexities compared to the kernel regime. For classification tasks, some focus has been on the XOR distribution(a.k.a. parities) [Wei et al., 2019]. Recent works have studied the problem of learning the $d$-dimensional XOR distribution using neural networks in both NTK and feature learning settings [Barak et al., 2022, Telgarsky, 2023, Taheri and Thrampoulidis, 2024, Glasgow, 2024]. Specifically, it has been shown that under NTK with a sufficiently small step size, a polynomially wide network requires $d^{2}$ GD steps and $d^{2}$ sample size. Some studies have achieved linear sample complexity for learning XOR [Bai and Lee, 2019, Glasgow, 2024, Telgarsky, 2023]; but these methods involve more computational effort compared to our results. The work most related to ours is [Glasgow, 2024], which demonstrated that with a particular Gaussian initialization, a ReLU network requires $\mathrm{poly}(\log(d))$ large SGD steps and $\mathrm{poly}(\log(d))$ neurons to learn XOR with linear sample complexity. In contrast, we show that by using a quadratic activation, learning this distribution requires only $\log(d)$ large steps with a constant-width network, while maintaining the same linear sample complexity.
@@ 2.1 Train and test loss bounds in deep nets
Consider the $L$-layer neural network with width $m$ where $\Phi(w,x):=\frac{1}{\sqrt{m}}W_{L+1}^{\top}(\frac{1}{\sqrt{m}}\sigma(W_{L}^{\top}\cdots\frac{1}{\sqrt{m}}\sigma(W_{1}^{\top}x)\cdots)$ and $\sigma$ is a $1$-smooth and $1$-Lipschitz activation function such that $\sigma(0)=0$. Moreover, let $\beta_{L}$ be a constant that only depends on $L$, let all parameters of the network be initialized as i.i.d. standard Gaussian and assume the step-size satisfies the condition of the descent lemma. Fix $T$ and assume the target weights vector $w^{\star}\in\mathbb{R}^{p}$ that obtain small training loss such that
@@
Moreover, assume for every $n$ samples from the data distribution there exists $w^{\star}$ satisfying Eqs. 2-3. Then, the expected generalization gap satisfies with high probability over the randomness of initialization,
@@
where the expectation is over the randomness in the training set denoted by $\mathcal{S}$ and $G_{0}$ is the Lipschitz parameter of network at initialization i.e., $\|\nabla\Phi(w_{0},\cdot)\|\leq G_{0}$.
@@ 2.1.1 Specializing to the NTK-separablity condition
In words, the above assumption implies that the features learned by the gradient at initialization can be linearly separated by some weights $w$. This assumption is commonly used in deep learning theory literature for studying classification tasks [Chen et al., 2021, Ji and Telgarsky, 2020].
@@
Consider the same setup as Theorem 2.1, let the loss function be logistic loss and let Assumption 1 hold. Define constant $B>0$ that bounds the model’s output at initialization i.e., $\forall i\in[n]:|\Phi(w_{0},x_{i})|<B$. Assume the width is large enough such that $m\geq\beta_{L}^{2}\left(\frac{2B+\log(1/{\epsilon})}{\gamma}\right)^{6L+4}.$ Then there exists $w^{\star}\in\mathbb{R}^{p}$ such that $F(w^{\star})\leq{\epsilon}$ and $w^{\star}$ lies at a bounded Euclidean distance from initialization such that $\|w^{\star}-w_{0}\|\leq\frac{1}{\gamma}(2B+\log(1/{\epsilon}))$.
@@
under the condition that $m=\Omega(\mathrm{poly}(\log(n))\cdot\beta_{L}/\gamma^{6L+4})$. We remark the bound does not have any explicit dependence on $m$. This in itself is not surprising as one intuitively expects the bound not to scale unfavorably with width. Although recent works have derived width independent generalization bounds for two-layer networks [Ji and Telgarsky, 2020, Telgarsky, 2023, Taheri and Thrampoulidis, 2024], we are not aware of any prior work proving width-independent bounds for learning multi-layer networks with GD. As discussed earlier in the introduction, the bound derived in the closely related work [Chen et al., 2021] scales unfavorably with $m$ as it grows at the rate $\sqrt{m/n}$. In fact, the authors refer to deriving width-independent bounds as an open problem in [Chen et al., 2021, Sec 3.1].
@@ 2.2 Overcoming NTK limitations for XOR dataset
Next, we show the previous NTK guarantees can be improved by using large step-sizes. We consider the stylized set-up of learning the $d-$dimensional XOR data distribution. Consider one-hidden layer network with quadratic activation where $x\in\mathbb{R}^{d},w_{i}\in\mathbb{R}^{d}$:
@@
Assume mini-batch SGD with batch-size $n\geq d\cdot\log^{14}(d)$ and step-size $\eta=m$. Then, after $T=\log(d)$ iterations the test accuracy satisfies with probability at least $1-e^{\log(m)-\log^{2}(d)}-e^{-\frac{m}{16}}-o_{d}(1)$ over the randomness of initialization and data sampling,
@@ 3 Numerical Results
Experiments on learning the XOR distribution with large step-size.
@@ B.2 Proof for Generalization loss
Lemma 5 (Lem. 6 in [Schliserman and Koren, 2022], Thm. 2 in [Lei and Ying, 2020]).
@@ E.4 Empirical Gradient for multiple GD steps
$w^{1}$ $=w^{0}-\eta\nabla F(w^{0}),$
$\mathbb{E}_{x}[w^{1}]$ $=w^{0}-\eta\mathbb{E}_{x}[\nabla F(w^{0})],$
$w^{1}$ $=\mathbb{E}_{x}[w^{1}]-\eta(\nabla F(w^{0})-\mathbb{E}_{x}[\nabla F(w^{0})]).$
@@
$w^{t+1}$ $=w^{t}-\eta\nabla F(w^{t}),$
$\mathbb{E}_{x}[w^{t+1}]$ $=\mathbb{E}_{x}[w^{t}]-\eta\mathbb{E}_{x}[\nabla F(\mathbb{E}_{x}[w^{t}])],$
$w^{t+1}$ $=\mathbb{E}_{x}[w^{t}]+(w^{t}-\mathbb{E}_{x}[w^{t}])-\eta(\nabla F(w^{t})-\mathbb{E}_{x}[\nabla F(\mathbb{E}_{x}[w^{t}])])-\eta\mathbb{E}_{x}[\nabla F(\mathbb{E}_{x}[w^{t}])].$
@@
$w^{t+1}-\mathbb{E}_{x}[w^{t+1}]$ $=-\eta\sum_{\tau=0}^{t}\nabla F(w^{\tau})-\mathbb{E}_{x}\left[\nabla F(\mathbb{E}_{x}[w^{\tau}])\right]$
$=-\eta\sum_{\tau=0}^{t}\underbrace{\nabla F(w^{\tau})-\nabla F(\mathbb{E}_{x}[w^{\tau}])}_{\text{Term I}}+\underbrace{\nabla F(\mathbb{E}_{x}[w^{\tau}])-\mathbb{E}_{x}[\nabla F(\mathbb{E}_{x}[w^{\tau}])]}_{\text{Term II}}.$
@@ E.5 Term II:
$-\nabla F_{i}(\mathbb{E}_{x}[w^{t}])=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w^{t}_{i}]\rangle x_{j}y_{j}$
@@
$-\nabla F_{i}^{k}(\mathbb{E}_{x}[w^{t}])$ $=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w_{i}^{t}]\rangle x_{j}(k)x_{j}(1)x_{j}(2)$
@@
$|\nabla F_{i}^{k}(\mathbb{E}_{x}[w^{t}])|$ $\lesssim\frac{\log(dT)\log(dm)}{nm}\left(\frac{2\sqrt{n}(|\alpha_{1}^{t}|+\alpha_{2}^{t})}{\sqrt{d}}+\sqrt{n}\log(d)\right)$
@@
$-\nabla F_{i}^{1}(\mathbb{E}_{x}[w^{t}])$ $=\frac{a_{i}}{nm}\sum_{j=1}^{n}\langle x_{j},\mathbb{E}_{x}[w_{i}^{t}]\rangle x_{j}(2)$
@@
$-\nabla F_{i}^{2}(\mathbb{E}_{x}[w^{t}])=\frac{a_{i}}{m}\mathbb{E}_{x}[w_{i}^{t}(1)]+O(\log(dm)\log(dT)\frac{|\alpha_{1}^{t}|+\alpha_{2}^{t}+\sqrt{d}\log(d)}{m\sqrt{nd}}).$ (50)
@@
$\left\|\nabla F_{i}(\mathbb{E}_{x}[w^{t}])-\mathbb{E}_{x}[\nabla F_{i}(\mathbb{E}_{x}[w^{t}])]\right\|\lesssim\log(dm)\log(dT)\frac{|\alpha_{1}^{t}|+\alpha_{2}^{t}+\sqrt{d}\log(d)}{m\sqrt{n}}.$
@@
$\forall i,t:\left\|\nabla F_{i}(\mathbb{E}_{x}[w^{t}])-\mathbb{E}_{x}[\nabla F_{i}(\mathbb{E}_{x}[w^{t}])]\right\|\lesssim\log(dm)\log(dT)\frac{|\alpha_{1}^{t}|+\alpha_{2}^{t}+\sqrt{d}\log(d)}{m\sqrt{n}},$